Suppose you performed a similar experiment using an unknown sample of carbonate. The pressure and volume were measured after the temperature stabilized at 21.5 ºC. Given the data in the table below, what is the stoichiometric ratio between HCl and CO2?

initial pressure (atm) 1.00

initial gas volume (mL) 240.0

volume of 1.00 M HCl solution added to fully consume the carbonate sample (mL) 22.0

final pressure (atm) 2.32

final gas volume (mL) 218.0
Answer #1

Moles of HCl added = conc. x volume = 1.00 M x ( 22/1000) L = 0.022 mol

final pressure = 2.32 atm

final gas volume = 218 mL = 0.218 L

Temperature = 21.5 C = 294.5 K

R = 0.0821 L.atm/mol.K

Moles of CO2 evolved = PV/RT = 2.32 x 0.218 / 0.0821 x 294.5 = 0.021 mol

the stoichiometric ratio between HCl and CO2 = 0.022 / 0.021 = 1.05